A 2.15 g sample of an unknown gas at 23 ∘C and 1.10 atm is stored in a 2.55 L

flask.

What is the density of the gas?

density:

g/L

What is the molar mass of the gas?

molar mass:
Answer #1

in Density = Mars Volume => Demity = 2.152 - 8 = 0.848/4